# Controlled Erasure as a Building Block for Universal Thermodynamically-Robust Superconducting Computing

### v1

@@ Abstract
Reducing the energy inefficiency of conventional CMOS-based computing devices—which rely on logically irreversible gates to process information—remains both a fundamental engineering challenge and a practical social challenge of increasing importance. In this work, we extend an alternative computing paradigm which utilizes distributions of microstates to store information in the metastable minima determined by an effective potential energy landscape. These minima serve as mesoscopic memory states which are manipulated by a dynamic landscape to perform information processing. Central to our results is the control erase (CE) protocol, which manipulates the landscape’s metastable minima to control whether information is preserved or erased. Importantly, successive executions of this protocol can implement a NAND gate—a logically-irreversible universal logic gate. We show how to practically implement this in a device created by two inductively coupled superconducting quantum interference devices (SQUIDs). We identify circuit parameter ranges that give rise to effective CEs and characterize the protocol’s robustness by comparing relevant protocol timescales. Due to their superconducting operational temperatures, performing optimized CEs with SQUIDs can serve as a platform for highly energy-efficient universal computation.
preprint: arXiv:XXXX.XXXXX
@@ I Introduction
Conventional classical computing systems harness irreversible logic gates—e.g., NAND gates—to process information. Irreversibility arises from erasing input information to create desired output states, coming at a minimum theoretical heat dissipation cost of $k_{B}T\ln 2$ [1] per erasure. During their operation, current conventional CMOS-based computational devices generate $\mathcal{O}(10^{4})$ times more heat than this minimum cost [2, 3, 4, 5]. On the social scale, energy consumption for computational purposes is projected to reach 20% of global energy demand by 2030 [6]. In light of this, it is time to investigate alternative strategies and substrates that give rise to energy-efficient universal computation.
One strategy implied in Landauer’s seminal 1961 result [1, 7, 8, 9, 10, 11, 12, 13, 14] focuses on manipulating metastable energy minima in an effective potential energy landscape. Coarse-graining the microstate phase-space surrounding the minima yields long-lived mesoscopic memory states that correspond to the logical 00s and $1$s for binary computation. From this perspective, a computation is implemented by controlling memory-state dynamics by changing the potential energy landscape. A computation is then a map between initial and final memory states. This “energy first” perspective of computation allows for careful analysis and prediction of a given logic gate’s performance and efficiency [1, 7, 11, 14].
Elaborating on this framework, Sec. II first describes the characteristics of a potential energy landscape capable of storing two bits of information. Sec. III introduces a family of control erase (CE) protocols—two-dimensional generalizations of single-bit erasure protocols. Sec. IV then presents a device constructed from two inductively coupled superconducting quantum interference devices (SQUIDs) [15, 16, 17, 18, 19, 20]. This device implements the required potential energy landscape and executes CE protocols by changing its circuit parameters. Sections V.1-V.3 characterize the effectiveness of the SQUID CE implementation by exploring circuit parameter ranges. Notably, with serial executions of CEs, a robust NAND gate can be performed, as detailed in Sec. V.4. Executing robust NAND gates with this device could lead to extremely low energy universal computations.
@@ II Physical Computing via Metastable Landscapes
We assume a potential energy landscape of interest supports several energy minima and an underlying physical substrate, which is connected to a thermal environment that introduces damping and noise into its dynamics. The minima are separated by energy barriers substantially larger than the thermal energy $k_{B}T$. The thermal environment quickly induces distributions of microstates to settle into local equilibria in the phase-space regions surrounding the minima. Noise, though, perturbs the microstates within their respective regions. However, it is unlikely to drive the microstates between these regions on a timescale that is exponential in energy barrier height. The energy barriers prevent mixing between these regions except on these very long timescales; as a result, the minima serve to support long-lived mesoscopic system states—metastable memory states. Manipulating them with the potential’s dynamics corresponds to information processing.
The following details physically-embedded computations via landscape control [1, 21, 22, 23, 13, 24]. Specifically, these computations are stochastic mappings from the set of initial to final memory states $\mathcal{M}$. Here, we introduce the control erase (CE) protocol over two input bits. One way of embedding 2-bit computations is to control a two-dimensional potential with a minimum in each quadrant. The memory states are distinguished using the $x$ and $y$ axes:
@@
Figure 1: Example 2-bit memory mesoscopic-state instantiation: Regions surrounding energy minima provide metastable information storage, whose bit value is assigned according to Eq. (1). Controlling this energy landscape manipulates the memory state’s dynamics—implementing information processing.
Fig. 1 illustrates a quadruple-well potential with an example memory-state instantiation chosen by the locations of the minima according to Eq. (1). The computational operations to be performed on the potential—i.e., the deterministic transformation of the potential that maps information from initial to final metastable memory states—are referred to as a physical protocol.
@@ III Control Erase Protocol
The NAND gate is a binary logical gate that takes in two input bits, meaning there are four possible input states: $00$, $01$, $10$, and $11$—and outputs one bit—either 00 or $1$. Specifically, the output state is $1$ for all input states except if both input bits are $1$, in which case the output is 00. Now, consider having no knowledge of a NAND gate’s input state and subsequently reading the output state to be $1$. Attempting to infer which specific two-bit input state led to this output state is impossible, as there are three possible input states that result in $1$. Microscopically, a state-space contraction occurred during the NAND gates operation, and this manifested as an irreversible erasure of information. However, if the output was 00, then the input is trivially known to be $11$. This is the underlying information processing task of the NAND gate: to preserve some input information, while erasing the rest. Thus, performing a NAND gate requires the ability to carry out controlled erasure protocols over the space of memory states.
@@
Figure 2: Illustrating the control erase (CE) protocol via the “dynamical skeleton” of the potential energy landscape in Fig. 1, which only shows the potential’s fixed points and corresponding flow fields. Here, stable minima (unstable saddle) fixed points are colored blue (red), while grey indicates a local maximum. This CE is executed with respect to the $y$ axis: The states contained on the negative $x$ half plane undergo an erasure of $\mathtt{Y}$ to the $1$ state, while those on the positive $x$ half plane do not change. This results in the third quadrant’s stored information being erased into the second quadrant. (Left) Initial potential energy landscape configuration. (Center) The stable and unstable fixed points in quadrant III approach each other just before annihilation. (Right) Just after fixed point annihilation in the third quadrant, the most important step of this example CE. If this landscape is held long enough, all information initially stored in quadrant III will be erased to quadrant II. Importantly, all other fixed points are maintained during this time. One completes the CE protocol cycle by bringing the potential back to its starting configuration.
@@
Erasure protocols in one-dimensional systems have been studied previously [1, 25, 11, 26, 13, 27, 28, 29]. The following extends this to an erasure protocol via the two-dimensional potential shown in Fig. 1. The additional freedom given by the second dimension permits control over what information is preserved and erased. Via that control, a NAND operation can be performed. To accomplish this, we introduce the control erase (CE) protocol. Fig. 2 illustrates a CE protocol via the potential’s skeleton, which views the potential from a dynamical systems perspective—i.e., its fixed points and flow fields in the microscopic state space. In this example, the negative $x$ half-plane executes an erasure protocol, while the positive $x$ half-plane remains unchanged.
@@
    First, choose which of the two input bits to operate on. For example, erasing information only along the $y$ axis requires operating only on the second bit so that $\mathtt{Y}^{\prime}\neq\mathtt{Y}$. This implies that the first input bit is preserved, i.e., $\mathtt{X}^{\prime}=\mathtt{X}$. However, this does not fully specify the pair of states that involved in the CE: $00$ and $01$, or $10$ and $11$, are both candidate pairs for a $\mathtt{Y}^{\prime}\neq\mathtt{Y}$ operation.
@@
    Next, specify which pair of two-bit memory states are involved in the erasure. Suppose we select states $10$ and $11$. Doing this controls the erasure operation on $\mathtt{X}$ such that it erases the information in $\mathtt{Y}$ only if $\mathtt{X}=1$. Otherwise, if $\mathtt{X}=0$, then the information stored in $\mathtt{Y}$ is preserved. This control will be denoted as $C_{\mathtt{X}}$.
@@
    Finally, we choose whether $\mathtt{Y}$ is erased to the 00 or the $1$ state. This is denoted as $E_{1}$, for example, if we erase to the state $1$;.
These three binary choices produce eight CE protocols of interest; which are detailed in Fig. 8. We denote the particular choices in the examples above with $\mathtt{Y}^{\prime}=C_{\mathtt{X}}E_{1}(\mathtt{X})$ and $\mathtt{X}^{\prime}=I_{\mathtt{X}}(\mathtt{X})$, where $I_{\mathtt{X}}$ indicates an identity operation of the $\mathtt{X}$ input bit , since the information from the first bit $\mathtt{X}$ is preserved. With this notation, and dropping the initial memory state dependence in the operations, the described CE protocol is written as: $(\mathtt{X}^{\prime},\mathtt{Y}^{\prime})=(I_{\mathtt{X}},C_{\mathtt{X}}E_{1})$.
Fig. 3 shows multiple representations of another CE: $(I_{\mathtt{X}},C_{\overline{\mathtt{X}}}E_{1})$. We are still erasing information stored in $\mathtt{Y}$, but instead of erasing the information if $\mathtt{X}=1$, the erasure happens if $\mathtt{X}=0$. In other words, our control is $\overline{\mathtt{X}}$ (NOT $\mathtt{X}$), rather than $\mathtt{X}$. Of particular importance is the arrow-based notation in Fig. 3, which will be employed when detailing the NAND gate in Sec. V.4.
Finally, to generalize the CE notation, first we know that one of the two output states $\rho^{\prime}\in\{\mathtt{X}^{\prime},\mathtt{Y}^{\prime}\}$ is created by the identity operation $\rho^{\prime}=I_{\rho}$, where $\rho\in\{\mathtt{X},\mathtt{Y}\}$. The other output state is created by $C_{\sigma}E_{A}$: Here, $\sigma\in\{\rho,\ \overline{\rho}\}$ indicates the control state for an erasure to the $A\in\{0,1\}$ state. Note that the overline $\overline{\rho}$ above $\rho$ indicates a negation of the $\rho$ state.
@@ IV Device Implementation
In this work, superconducting quantum interference devices (SQUIDs) are employed as a physical platform for performing CE protocols. We first briefly review relevant superconducting device literature. Then, we specify the device of interest—two inductively coupled SQUIDs (SQUIDs), shown in Fig. 4—whose potential energy landscape can dynamically execute CE protocols.
The quantum flux parametron (QFP) was invented by Goto in 1985 [30, 15], whose primary focus was for energy-efficient classical computing applications in the superconducting regime. With the incentive of optimizing circuit parameters and operating on slower computational timescales to aid in energy efficiency, Takeuchi et al. introduced the adiabatic QFP [31, 4]. In contrast to our work, circuit parameters are not optimized via algorithms for thermodynamic performance, but are instead found by utilizing bifurcation theory. With a similar design construction as the QFP, but having the same equations of motion as the QFP, in 1989 Han et al. introduced the variable $\beta$ radio frequency SQUID [16, 18, 17]. This device was later named the compound Josephson junction radio frequency SQUID (CJJ rf SQUID) [32, 19, 33]. Its principle use was for investigating macroscopic quantum phenomena (MQP). The goal of the present effort—demonstrating universal classical information and storage—is not shared with the CJJ rf SQUID literature.
This said, the circuit parameter values from these areas of literature is used in Sec. V.3 for evaluating the performance of the device of interest when performing CE protocols. Since QFPs and CJJ rf SQUIDs are both SQUIDs, we refer to the device in Fig. 4 as two inductively coupled SQUIDs. Ref. [20] introduced and detailed the device in Fig. 4. Importantly, it supports a potential equivalent to that of Fig. 1. To understand how it executes computations, the following gives a synopsis of the device’s potential energy landscape created. Physically-grounded approximations and assumptions for executing a CE protocol are then detailed.
@@
Fig. 4 illustrates a device consisting of two inductively coupled SQUIDs. In this circuit, $L_{i}$ [$l_{j}$] indicate the radio frequency [direct current] SQUID inductances, while $J_{j}$ indicate Josephson junctions, all for which $i=1,2$, and $j=1,2,3,4$. This device generates the following potential energy surface:
@@
Here, $U_{0}=(\Phi_{0}/2\pi)^{2}/L_{\alpha}$, where $\Phi_{0}$ is the flux quantum, $\varphi_{r}=(2\pi\phi_{r})/\Phi_{0}$ is the reduced flux variable of the $r$th flux variable, and $L_{\alpha}=\alpha L$ for which $\alpha=1-M^{2}/L^{2}$ with $M$ being a tunable mutual inductance between the two SQUIDs. Next, $L\coloneqq L_{1}\approx L_{2}$, $\gamma_{i}=L_{\alpha}/2l_{i}$ in which $l_{1(2)}\coloneqq l_{1(3)}\approx l_{2(4)}$, $\beta_{n}=2\pi L_{\alpha}(I_{\mathrm{c}2(4)}+I_{\mathrm{c}1(3)})/\Phi_{0}$, and $\delta\beta_{n}=2\pi L_{\alpha}(I_{\mathrm{c}2(4)}-I_{\mathrm{c}1(3)})/\Phi_{0}$, where each $I_{\mathrm{c}j}$ corresponds to the critical current $I_{\mathrm{c}}$ of the $j$th Josephson junction, with $n=1,2$. Lastly, $\mu=M/L$. Here, $M$ is treated as a tunable constant: While its experimental implementation—a SQUID coupler—are discussed in Refs. [34, 32, 33], the coupler’s dynamics will not be addressed here.
Note that Eq. (2) contains four degrees of freedom, namely $\varphi_{i}$ and $\varphi_{i\mathrm{dc}}$, where $i=1,2$. This type of compound SQUID is generally constructed with $\gamma\gg 1$ [18, 35]. This implies that any changes in $\varphi_{i\mathrm{xdc}}$ are rapidly observed in $\varphi_{i\mathrm{dc}}$, so we assume $\varphi_{i\mathrm{dc}}=\varphi_{i\mathrm{xdc}}$. Applying this approximation to Eq. (2) projects the potential onto two dimensions, providing the potential energy surface shown in Fig. 1 in the $\varphi_{1}$$\varphi_{2}$ plane.
Next, we assume that the fabrication consistency of the JJ elements is such that $\delta\beta_{1}=\delta\beta_{2}=0$. In practice, $\delta\beta_{1}\neq\delta\beta_{2}\neq 0$, which results in a slight offset in the minima’s locations [27]. This induced asymmetry in a real device can be at least partially compensated by external flux parameters if necessary. The $\delta\beta=0$ assumption is fairly common, and is made in order to streamline the process of analyzing the device’s equation of motion. As a similar assumption, we let $\beta\coloneqq\beta_{1}=\beta_{2}$ and, furthermore, $I_{\mathrm{c}}\coloneqq I_{\mathrm{c}1}=I_{\mathrm{c}2}=I_{\mathrm{c}3}=I_{\mathrm{c}4}$.
Finally, assuming small-valued coupling ratios, we drop terms that are quadratic in $\mu$, which yields $\alpha\rightarrow 1$ and $L_{\alpha}\rightarrow L$.
@@
Eq. (3) describes a dimensionless potential $U^{\prime}$ with two degrees of freedom; this will be frequently referred to for the remainder of this work. We manipulate this potential energy landscape with the following circuit parameters for $i=1,2$: (i) $\varphi_{i\mathrm{x}}$ tilts the potential with respect to the $i$th axis, (ii) $\varphi_{i\mathrm{xdc}}$ provides an in-situ barrier control on the $i$th axis, and (iii) $\mu$ supplies a coupling interaction between the two circuits, which biases the potential to favor wells that lie on one of the two diagonals in the $\varphi_{1}$$\varphi_{2}$ plane.
V Contol Erase Protocol Implementation
The approximations in Sec. IV yield the potential given in Eq. (3), and is displayed in Fig. 1. Manipulating the aforementioned external circuit parameters implements any of the eight possible control erase (CE) protocols shown in Fig. 8. Using the CE $(I_{\mathtt{X}},C_{\overline{\mathtt{X}}}E_{1})$ from Figs. 2 and 3 as an example, we detail how to find circuit parameter values that give rise to “effective” protocols in Sec. V.1. We qualitatively show this by exploring circuit parameter values that give rise to ranges for which this particular CE can be executed. Furthermore, we quantify this protocol’s effectiveness—i.e., if it is highly robustness—in Sec. V.3. Sec. V.2 generalizes to determine which circuit parameters will be used to carry out any CE protocol in Fig. 8. With this, Sec. V.4 demonstrates the physical execution of an effective NAND gate.

### v2

@@ Abstract
Reducing the energy inefficiency of conventional CMOS-based computing devices—which rely on logically irreversible gates to process information—remains both a fundamental engineering challenge and a practical social challenge of increasing importance. We extend an alternative computing paradigm that manipulates microstate distributions to store information in the metastable minima determined by an effective potential energy landscape. These minima serve as mesoscopic memories that are manipulated by a dynamic landscape to perform information processing. Central to our results is the control erase (CE) protocol that controls the landscape’s metastable minima to determine whether information is preserved or erased. Importantly, successive protocol executions can implement a NAND gate—a logically-irreversible universal logic gate. We show how to practically implement this in a device created by two inductively-coupled superconducting quantum interference devices (SQUIDs). We identify circuit parameter ranges that give rise to effective CEs and establish the device’s robustness against logical errors. These SQUID-based logical devices are capable of operating above GHz frequencies and at the $k_{B}T$ energy scale. Due to this, optimized devices and associated protocols provide a universal-computation substrate that is both computationally fast and energy efficient.
preprint: arXiv:2406.12153
@@ I Introduction
Conventional classical computing systems harness irreversible logic gates—e.g., NAND gates—to process information. Irreversibility arises from erasing input information to create desired output states, coming at a minimum theoretical heat dissipation cost of $k_{B}T\ln 2$ [1] per erasure. During their operation, current conventional CMOS-based computational devices generate $\mathcal{O}(10^{4})$ times more heat than this minimum cost [2, 3, 4, 5]. On the social scale, energy consumption for computational purposes is projected to reach 20% of global energy demand by 2030 [6]. In light of this, it is time to investigate alternative strategies and substrates that give rise to markedly-more energy-efficient universal computation.
One strategy implied in Landauer’s seminal 1961 result [1, 7, 8, 9, 10, 11, 12, 13, 14] focuses on manipulating metastable energy minima in an effective potential energy landscape. Coarse-graining the microstate phase-space surrounding the minima yields long-lived mesoscopic memory states that correspond to the logical 00s and $1$s for binary computation. From this perspective, a computation is implemented by controlling memory-state dynamics through changes in the landscape. Altogether, a computation is then a map between initial and final memory states. This “energy first” perspective of computation allows for careful analysis and prediction of a given logic gate’s performance and efficiency [1, 7, 11, 14].
Elaborating on this framework, Sec. II first describes a family of potential energy landscapes capable of storing two bits of information. Section III introduces a family of control erase (CE) protocols—two-dimensional generalizations of single-bit erasure protocols. Section IV then presents a device constructed from two inductively coupled superconducting quantum interference devices (SQUIDs) [15, 16, 17, 18, 19, 20]. We demonstrate that this device implements the required family of potential energy landscapes and executes CE protocols by changing its circuit parameters. Sections V.1-V.3 explore the effectiveness of the SQUID CE implementation as a function of circuit parameters. Notably, serial executions of CEs produce a robust NAND gate, as detailed in Sec. V.4. Executing robust NAND gates with this device supports extremely low-energy universal computation.
@@ II Physical Computing via Metastable Landscapes
Consider a potential energy landscape that exhibits several energy minima supported via an underlying physical substrate which itself is connected to a thermal environment that introduces damping and noise into its dynamics. The minima are separated by energy barriers, whose heights are substantially larger than the thermal energy $k_{B}T$. The thermal environment quickly induces distributions of microstates to settle into local equilibria in the phase-space regions surrounding these wells. Noise perturbs the microstates within their respective regions; however, it is unlikely to drive the microstates between these regions on a timescale that scales exponentialy with the energy barrier height. The energy barriers prevent mixing between these regions except on these very long timescales. As a result, the minima serve to support long-lived mesoscopic system states—metastable memory states. Manipulating them with the potential’s dynamics corresponds to information processing.
The following details physically-embedded computations via landscape control in this setting [1, 21, 22, 23, 13, 24]. Specifically, these computations are stochastic mappings from the set of initial to final memory states $\mathcal{M}$. Here, we introduce the control erase (CE) protocol over two input bits. One way of embedding 2-bit computations is to control a two-dimensional potential with a minimum in each quadrant. The memory states are distinguished using the $x$ and $y$ axes:
@@
Figure 1: Example 2-bit memory mesoscopic-state instantiation in a four well potential. Regions surrounding energy minima provide metastable information storage, whose bit value is assigned according to Eq. (1). Controlling this energy landscape manipulates the memory state’s dynamics, thereby implementing information processing.
Figure 1 illustrates a quadruple-well potential with an example memory-state instantiation chosen by the locations of the minima according to Eq. (1). The computational operations to be performed on the potential—i.e., the deterministic transformation of the potential that maps information from initial to final metastable memory states—is referred to as a physical protocol.
@@ III Control Erase Protocol
The NAND gate is a binary logical gate that takes in two input bits, meaning there are four possible input states—$00$, $01$, $10$, and $11$—and one output bit—either 00 or $1$. Specifically, the output state is $1$ for all input states except if both input bits are $1$, in which case the output is 00.
Consider reading the output state to be $1$ with no knowledge of a NAND gate’s input state. Since there are three possible inputs that can produce this single output state, inferring which specific two-bit input state led to the output bit $1$ is impossible. The logical impossibility manifests as a physically-irreversible erasure of information associated with phase-space contraction at the microscopic scale. In contrast, if the output was 00, then the input is trivially known to be $11$ and no phase-space contraction is entailed.
The above illustrates the NAND’s underlying information processing task: Preserve some input information, while erasing the rest. In this way, performing a NAND gate requires the ability to carry out controlled erasure protocols over the space of memory states.
@@
Figure 2: Control erase (CE) protocol illustrated via the dynamical skeleton of the potential energy landscape in Fig. 1, showing only the potential’s fixed points and corresponding local flow fields. Stable minima (unstable saddle) fixed points are colored blue (red), while grey indicates a local maximum. This CE is executed with respect to the $y$ axis: The states contained on the negative $x$ half plane undergo an erasure of $\mathtt{Y}$ to the $1$ state, while those on the positive $x$ half plane do not change. This results in the third quadrant’s stored information being erased into the second quadrant. (Left) Initial potential energy landscape configuration. (Center) The stable and unstable fixed points in quadrant III approach each other just before annihilation. (Right) Immediately after fixed point annihilation in the third quadrant. If this landscape is held long enough, all information initially stored in quadrant III will be erased to quadrant II. Importantly, all other fixed points are maintained during this time. One completes the CE protocol cycle by bringing the potential back to its starting configuration.
@@
Erasure protocols in one-dimensional systems have been studied previously [1, 25, 11, 26, 13, 27, 28, 29]. The following extends this to an erasure protocol via the two-dimensional potential shown in Fig. 1. The additional freedom given by the second dimension permits control over what information is preserved and erased. Via that control, a NAND operation can be performed. To accomplish this, we introduce the control erase (CE) protocol. Figure 2 illustrates a CE protocol via the potential’s skeleton, viewing the potential from a dynamical systems perspective—i.e., via its fixed points and local flow fields in the microscopic state space. In this example, the negative $x$ half-plane executes an erasure protocol, while the positive $x$ half-plane remains unchanged.
@@
    First, choose which of the two input bits to operate on. For example, erasing information along the $y$ axis requires operating only on the second bit so that $\mathtt{Y}^{\prime}\neq\mathtt{Y}$. This implies that the first input bit is preserved, i.e., $\mathtt{X}^{\prime}=\mathtt{X}$. However, this does not fully specify the pair of states that involved in the CE: $00$ and $01$, or $10$ and $11$, are both candidate pairs for a $\mathtt{Y}^{\prime}\neq\mathtt{Y}$ operation.
@@
    Next, specify which pair of two-bit memory states are involved in the erasure. Suppose we select the states $10$ and $11$: Doing this controls the erasure operation on $\mathtt{X}$ such that information is erased in $\mathtt{Y}$ only if $\mathtt{X}=1$. Otherwise, if $\mathtt{X}=0$, the information stored in $\mathtt{Y}$ is preserved. We denote this choice as a function $CE(\mathtt{X},\mathtt{Y})$ that is an identity on $\mathtt{Y}$ if $\mathtt{X}=0$ and an erasure of $\mathtt{Y}$ if $\mathtt{X}=1$.
@@
    However, we also need to choose whether $\mathtt{Y}$ is erased to the 00 or the $1$ state. For example, if we erase to the state $1$, we augment our notation to be $CE_{1}(\mathtt{X},\mathtt{Y})$.
These three binary choices produce the eight CE protocols of interest shown in Fig. 8. All said, the choices above give us $(\mathtt{X}^{\prime},\mathtt{Y}^{\prime})=(I(\mathtt{X}),CE_{1}(\mathtt{X},\mathtt{Y}))$, where $I(\mathtt{X})$ indicates an identity operation on the $\mathtt{X}$ input bit. For legibility, we introduce a condensed notation that suppresses the functional dependencies: $(\mathtt{X}^{\prime},\mathtt{Y}^{\prime})=(I_{\mathtt{X}},C_{\mathtt{X}}E_{1})$. The following relies on this notation.
Figure 3 shows multiple representations of another CE: $(I_{\mathtt{X}},C_{\overline{\mathtt{X}}}E_{1})$. This also erases information stored in $\mathtt{Y}$, but instead of erasing the information if $\mathtt{X}=1$, the erasure happens if $\mathtt{X}=0$. In other words, our control is $\overline{\mathtt{X}}$ (NOT $\mathtt{X}$), rather than $\mathtt{X}$. Of particular importance is the arrow-based notation in Fig. 3, which is employed when detailing the NAND gate in Sec. V.4.
Finally, to generalize the CE notation, first we know that one of the two output states $\rho^{\prime}\in\{\mathtt{X}^{\prime},\mathtt{Y}^{\prime}\}$ is created by the identity operation $\rho^{\prime}=I_{\rho}$, where $\rho\in\{\mathtt{X},\mathtt{Y}\}$. The other output state is created by $C_{\sigma}E_{A}$. Here, $\sigma\in\{\rho,\ \overline{\rho}\}$ indicates the control state for an erasure to the $A\in\{0,1\}$ state. Note that the overline $\overline{\rho}$ above $\rho$ indicates a negation of the $\rho$ state. The eight CEs of interest are shown in Fig. 8 in App. C.
@@ IV Device Implementation
Under the current computing paradigm, superconducting quantum interference devices (SQUIDs) are employed only as a physical platform for performing CE protocols. (To emphasize, other than zero-resistance supercurrents and the Josephson nonlinearity, no quantum-mechanical phenomena are used for information processing. And, low-temperate operation is relied on only insofar as it freezes out irrelevant thermal degrees of freedom.) We first briefly review related superconducting device physics. Then, we specify the device of interest—two inductively coupled SQUIDs, shown in Fig. 4—whose potential energy landscape can dynamically execute CE protocols.
Historically, the quantum flux parametron (QFP) was invented by Goto in 1985 [30, 15] for classical computing applications on an energy-efficient superconducting platform. With the incentive of optimizing circuit parameters and operating on slower computational timescales to aid energy efficiency, Takeuchi et al introduced the adiabatic QFP [31, 4]. In contrast to our focus here, their circuit parameters were not optimized for thermodynamic performance, but were instead were determined via bifurcation theory. With a similar design construction as the QFP, but having the same equations of motion as the QFP, in 1989 Han et al introduced the variable $\beta$ radio frequency SQUID [16, 18, 17]. This device was later named the compound Josephson junction radio frequency SQUID (CJJ rf SQUID) [32, 19, 33]. Its principle use was for investigating macroscopic quantum phenomena (MQP), not classical information processing. As such, our present goal—demonstrating universal classical information and storage—is not shared with the CJJ rf SQUID literature.
This said, the circuit parameter values used in these earlier devices happen to be useful in Sec. V.3 for evaluating the performance of our device of interest when performing CE protocols. Since QFPs and CJJ rf SQUIDs are both SQUIDs, we refer to the device in Fig. 4 as two inductively coupled SQUIDs. Reference [20] introduced and detailed the device in Fig. 4. Importantly, it supports a potential equivalent to that of Fig. 1. To understand how it executes computations, the following gives a synopsis of the device’s potential energy surface. Physically-grounded approximations and assumptions for executing a CE protocol are then detailed.
@@
Figure 4 illustrates a device consisting of two inductively coupled SQUIDs. In this circuit, $L_{i}$ [$l_{j}$] indicate the radio frequency [direct current] SQUID inductances, while $J_{j}$ indicate Josephson junctions, all for which $i=1,2$, and $j=1,2,3,4$. This device generates the following potential energy surface:
@@
Here, $\varphi_{r}=(2\pi\phi_{r})/\Phi_{0}$ is the reduced flux variable of the $r$th flux variable, while $\Phi_{0}$ is the flux quantum. We subsequently take $L\coloneqq L_{1}=L_{2}$ and $l_{1(2)}\coloneqq l_{1(3)}=l_{2(4)}$. Next, $L_{\alpha}=\alpha L$ for which $\alpha=1-\mu^{2}$ such that $\mu=M/L$. $M$ is a tunable mutual inductance constant: While its experimental implementation—a SQUID coupler between the two SQUIDs—are discussed in Refs. [34, 32, 33], the coupler’s dynamics will not be addressed here. With this, $U_{0}=(\Phi_{0}/2\pi)^{2}/L_{\alpha}$ and $\gamma_{1(2)}=L_{\alpha}/2l_{1(2)}$. Last, $\beta_{1(2)}=2\pi L_{\alpha}(I_{\mathrm{c}2(4)}+I_{\mathrm{c}1(3)})/\Phi_{0}$, and $\delta\beta_{1(2)}=2\pi L_{\alpha}(I_{\mathrm{c}2(4)}-I_{\mathrm{c}1(3)})/\Phi_{0}$, where each $I_{\mathrm{c}j}$ corresponds to the critical current $I_{\mathrm{c}}$ of the $j$th Josephson junction.
Note that Eq. (2) contains four degrees of freedom, namely $\varphi_{i}$ and $\varphi_{i\mathrm{dc}}$, where $i=1,2$. This type of compound SQUID is generally constructed with $\gamma_{i}\gg 1$ [18, 35]. This implies that any changes in $\varphi_{i\mathrm{xdc}}$ are rapidly observed in $\varphi_{i\mathrm{dc}}$, so we assume $\varphi_{i\mathrm{dc}}=\varphi_{i\mathrm{xdc}}$. Applying this approximation to Eq. (2) projects the potential onto two dimensions, providing the potential energy surface shown in Fig. 1 in the $\varphi_{1}$$\varphi_{2}$ plane.
Next, we assume that the fabrication consistency of the JJ elements is such that $\delta\beta_{1}=\delta\beta_{2}=0$: This assumption is fairly common, as it streamlines the process of analyzing the device’s equation of motion. In practice, $\delta\beta_{1}\neq\delta\beta_{2}\neq 0$, which results in a slight offset in the wells’ locations [27]. This induced asymmetry in a real device can be at least partially compensated by external flux parameters if necessary. Along similar lines, we let $\beta\coloneqq\beta_{1}=\beta_{2}$ as well as $I_{\mathrm{c}}\coloneqq I_{\mathrm{c}1}=I_{\mathrm{c}2}=I_{\mathrm{c}3}=I_{\mathrm{c}4}$. Finally, assuming small-valued coupling ratios, we drop terms that are quadratic in $\mu$, which sends $\alpha\rightarrow 1$ and $L_{\alpha}\rightarrow L$.
@@
Equation (3) describes a dimensionless potential $U^{\prime}$ with two degrees of freedom; this will be frequently referred to in the remainder. We manipulate the potential energy landscape with the following circuit parameters for $i=1,2$: (i) $\varphi_{i\mathrm{x}}$ tilts the potential with respect to the $i$th axis, (ii) $\varphi_{i\mathrm{xdc}}$ provides an in-situ barrier control on the $i$th axis, and (iii) $\mu$ supplies a coupling interaction between the two circuits, which biases the potential to favor wells that lie on one of the two diagonals in the $\varphi_{1}$$\varphi_{2}$ plane.
V Control Erase Protocol Implementation
The approximations in Sec. IV yield the potential given in Eq. (3) and displayed in Fig. 1. Manipulating the aforementioned external circuit parameters implements any of the eight possible control erase (CE) protocols shown in Fig. 8. Using the CE $(I_{\mathtt{X}},C_{\overline{\mathtt{X}}}E_{1})$ from Figs. 2 and 3 as an example, we detail how to find circuit parameter values that give rise to “effective” protocols in Sec. V.1. We qualitatively show this by exploring circuit parameter values that give rise to ranges for which this particular CE can be executed. Furthermore, we quantify this protocol’s effectiveness—i.e., if it is highly robust—in Sec. V.3. Section V.2 generalizes to determine which circuit parameters will be used to carry out any CE protocol in Fig. 8. With this, Sec. V.4 demonstrates the physical execution of an effective NAND gate.